OpenStudy (anonymous):

How many solutions are there to the equation 5x+48+7x=12(x+4)

OpenStudy (anonymous):

How many solutions are there to the equation 5x+48+7x=12(x+4)

OpenStudy (anonymous):

Any number B/C it turns into 12x + 48= 12x + 48